Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

Tel: 0703 593081

email: dbc@uk.ac.soton.ecs

June 8, 1994

Geoffrey C. Fox Concurrency: Practice and Experience Syracuse University Northeast Parallel Architectures Center 111 College Place Syracuse, New York 13244-4100

Dear Prof. Fox

I apologise for the delay in refereeing the paper

“Performance evaluation of load balancing algorithms for parallel singlephase iterative PDE solvers.”

by Nikos Chisochoides, Nashat Mansour and Geoffrey Fox. In case it is not too late, please find enclosed my opinions.

Yours sincerely,